OpenStudy (anonymous):

Help please!!!! Arthur has decided to start saving for a new computer. His money is currently in a piggy bank at home, modeled by the function s(x) = 85. He was told that he could do the laundry for the house and his allowance would be a(x) = 10(x − 1), where x is measured in weeks. Explain to Arthur how he can create a function that combines the two, and describe any simplification that can be done.

OpenStudy (anonymous):

@phi

OpenStudy (amorfide):

s(x)=85 piggy bank a(x)=10(x-1) allowance he currently has the money in his piggy bank which is 85 then he gains 10(x-1) for every week that x is, so you know that on the first week he is doing laundry he has the allowance a(1)=0 but you already have 85 in your piggy bank so your new function will give you 0+85 meaning 85 is your starting case so you always have 85 from the start therefore you get a new function s(x)+a(X) which we will label f(x) so f(x)=s(x)+a(x) f(x)=85+ 10(x-1) @texaschic101 confirm? ( I suck at this )

OpenStudy (anonymous):

It right to me thank you so much !!!

OpenStudy (amorfide):

no problem, always happy to help

OpenStudy (anonymous):

TY!!!

OpenStudy (anonymous):

THIS REALLY HELPED!!